OpenStudy (thedj4jc):

What is the excluded value? 2x/1-2x A. 0 B. 1/2 C. 1

OpenStudy (thedj4jc):

And... What are the excluded values? \[\frac{ -7 }{ x^2-2x-15 }\] 0 and 15 3 and -5 -3 and 5

OpenStudy (anonymous):

find the value of x for which denominator is zero

OpenStudy (thedj4jc):

1. B But I can't find the answer for the second problem. I have... A = -45 B = 4 C = -34 @surjithayer

OpenStudy (anonymous):

\[x ^{2}-2x-15=0\] 2=5-3 5*-3=-15 \[x^2-(5-3)x-15=0\] \[x^2-5x+3x-15=0 \] make factors and find values of x

OpenStudy (thedj4jc):

I'm lost

OpenStudy (anonymous):

x=-3 or x=5 put in equation will=0 can't divide by zero